Given P(A) = 0.32, P(B) = 0.48, P(A or B) = 0.36, what is P(A and B)?
Found 2 solutions by Theo, greenestamps:
Answer by Theo(13342)

p(a or b) = p(a) + p(b) - p(a and b)
you get:
.36 = .32 + .48 - p(a and b)
solve for p(a and b) to get:
p(a and b) = .32 + .48 - .36 = .44
you get:
p(a or b) = p(a) + p(b) - p(a and b) becomes .36 = .32 + .48 - .44 = .36, confirming the arithmetic is correct.
here's a referrence.
https://www.statology.org/probability-of-a-or-b/

Answer by greenestamps(13327)



The given information is impossible. P(A or B) can't be less than P(B).

ANSWER: None -- the problem is faulty